# A Family of Self-Dual Quasicrystals with Critical Phases

### v2

@@ Abstract
We propose a general framework for constructing self-dual one-dimensional quasiperiodic lattice models with arbitrary-range hoppings and multifractal behaviors. Our framework generates a broad spectrum of quasicrystals, ranging from the off-diagonal Aubry-André-Harper models on one end, to those with long-range power-law hoppings on another. Focusing on models with off-diagonal quasiperiodic hoppings with power-law decay, we exploit the fact that, when the self-dual condition is satisfied, the system must be in the critical state with multifractal properties. This enables the engineering of models with competing extended, critical, and localized phases, with rich mobility edges in between. As an outstanding example, we show that a limiting case of our family of self-dual quasicrystals can be implemented using Rydberg-atom arrays. Our work offers a systematic route toward self-duality and critical phases, and would facilitate their experimental simulation.
@@
Introduction.— Anderson localization is a fundamental quantum phenomenon in which disorder leads to the exponential localization of eigenstates, thereby inhibiting wave diffusion [1, 2, 3, 4, 5, 6]. The associated location-delocalization transition generally depends on the spatial dimensions, for instance, an infinitesimally weak disorder is sufficient to induce Anderson localization in one dimension [7, 8, 9]. But Anderson localization also arises in quasiperiodic models, where, in a one dimensional quasicrystal for instance, the localization only occurs beyond a critical threshold of the quasiperiodic potential. This is exemplified by the well-known Aubry-André-Harper (AAH) model [10], characterized by the eigen equation $t\left(\psi_{n+1}+\psi_{n-1}\right)+V\cos(2\pi\tau n)\psi_{n}=E\psi_{n}$, where $\psi_{n}$ is the wavefunction component on site $n$, $t$ is the nearest-neighbor hopping strength, $V$ is the amplitude of the quasiperiodic potential, $\tau$ is an irrational number, and $E$ is the eigenenergy. Crucially, the AAH model exhibits self-duality, such that it Fourier-transforms onto itself but with switched parameters $V\leftrightarrow 2t$. The self-dual point $V=2t$ thus marks a sharp transition between the extended (for $V<2t$) and localized (for $V>2t$) phases. Remarkably, at the self-dual point, the system becomes critically localized [12, 13, 14, 15, 16, 17, 11, 18, 19, 20, 21].
@@
Further engineering of the AAH model can have two important consequences. First, with the addition of off-diagonal quasiperiodic potentials, the critical point can be expanded into a critical phase occupying a broad region on the phase diagram. These critical states are extended yet nonergodic, exhibiting unique spectral statistics [22, 23, 24], multifractal properties [25, 26, 27], and unconventional dynamics [28, 29, 30]. Second, by introducing long-range hopping terms [31] or fine-tuned on-site quasiperiodic potentials [32], mobility edges emerge in the eigenspectum, where critical energies separate extended and localized states [3, 5, 31, 32, 34, 35, 33, 36]. These possibilities have sparked widespread interest in quasiperiodic systems, including theoretical studies on periodically (or quasiperiodically) driven systems [37, 38, 39, 40, 41, 42] and non-Hermitian quasicrystals [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], as well as experimental implementations using ultracold atoms [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], photonic crystals [65, 66, 67, 68, 69, 70, 71, 72], optical cavities [73, 74], and superconducting circuits [75, 76]. Despite these activities, the general mechanisms responsible for the emergence of critical states remain an open question, and the coexistence (in the same phase diagram) of extended, critical, and localized phases has so far been observed only in a handful of fine-tuned models. While it is tempting to address these issues from the perspective of self-duality, only a limited number of quasiperiodic models are known to be self-dual.
@@
In this work, we propose a general framework for constructing self-dual one-dimensional quasiperiodic models with arbitrary-range hopping and critical behaviors. Similar to the AAH example, with our scheme, a model constructed with the parameters $(a,b)$ is dual to the same model but parameterized by $(b,a)$. This naturally leads to the self-dual condition $a=b$, where the system must be in the critical state. The family of quasicrystals so generated spans a wide spectrum, with the (generalized) AAH models on one end, and those with long-range hoppings on another. Inspired by the recent experimental progress of Rydberg-atom arrays, we focus on models exhibiting off-diagonal quasiperiodic hoppings with power-law decay. Consistent with our expectation, the system is in the critical phase close to the self-dual point. The overall phase diagram is symmetric with respect to the self-dual condition $a=b$, featuring rich mobility edges in between the extended, critical, and localized phases. We then demonstrate that the model parameterized by $(3,+\infty)$, which features both extended and critical phases, can be experimentally realized using Rydberg atoms with dipole-dipole interactions. Our work provides a general scheme for engineering self-dual quasicrystals, as well as a route toward critical phases from self-duality considerations.
@@
where the irrational number $\tau=({\sqrt{5}-1})/{2}$. Performing the dual transformation
@@
where $N$ represents the total number of sites, the Hamiltonian in the dual space reads [77]
@@
Equations (1) and (4) establish that $\hat{H}(a,b)$ and $\hat{H}(b,a)$ are dual to each other, such that $a=b$ defines the self-dual condition. It follows that, if the $j$th eigenstate of $\hat{H}(a,b)$, denoted as $\psi^{(j)}$, is localized, then the $j$th eigenstate of $\hat{H}(b,a)$, denoted as $\tilde{\psi}^{(j)}$, must be extended, and vice versa. Furthermore, if $\psi^{(j)}$ is critical, then $\tilde{\psi}^{(j)}$ is also critical, such that all eigenstates of $\hat{H}(a,a)$ are critical, being at the self-dual point $a=b$. Note that, in the following, we index the eigenstates according to their eigenenergies in ascending order.
Here two remarks are in order. First, Hamiltonian $\hat{H}(a,b)$ is determined by the functions $\{f_{s}(a)\}_{s=1,2,...,+\infty}$, which are arbitrary, and subject only to the constraint that the summation in Eq. (2) should converge. Second, such a flexibility allows us to construct a broad class of self-dual quasicrystals with multifractal properties [77].
@@
Quasicrystals with power-law hoppings.— Inspired by the recent progress in Rydberg-atom arrays with long-range interactions, we now focus on a class of qusiperiodic models with power-law hopping terms. Specifically, we consider $f_{s}(a)=\frac{1}{s^{a}}$ for $s\leq d$, and $f_{s}(a)=0$ for $s>d$. In this case, we have
@@
where $d$ denotes the maximum hopping range, while $a$ and $b$ respectively characterize the exponent of the power-law decay in hopping, and the amplitude modulation.
To corroborate the self-duality and localization properties of the model, we study the fractal dimension (FD) of $\hat{H}_{p}(a,b)$. The fractal dimension is given by
@@
As a convenient sampling of the parameter space, we impose the constraint $a+b=4$, such that the self-dual condition is satisfied for $a=2$. In Fig. 1(a)(b), we show the computed FD as functions of $a$ and the eigenstate index $j/N$ for different hopping ranges $d$. The resulting phase diagram exhibits a rich structure, featuring all three phases: the extended (red), the critical (green), and the localized (blue). Consistent with the self-dual properties of the model, the phase diagram is symmetric: regions of the extended (localized) phase to the left of $a=2$ are symmetric to those of the localized (extended) phase to the right, while regions of the critical phase are symmetric with respect to $a=2$. With increasing hopping range $d$, the stability region of the critical phase shrinks and becomes fragmented.
@@
In Fig. 1(c), we show the fractal dimension as a function of the eigenstate index $j/N$ for different system sizes. In the localized regions (marked I and IV), FD tends toward zero with increasing system size, while in the extended region (marked III), FD increases toward unity. In contrast, in the critical region (marked II), FD remains between 0 and 1, showing minimal dependence on the system size.
@@
Alternatively, the three phases can be distinguished by analyzing the level statistics. Denoting the ordered eigenenergies as $\{E_{j}\}_{j=1,2,...N}$, we define the even-odd eigen-level spacing as $s_{2j}^{e-o}=E_{2j}-E_{2j-1}$, and the odd-even eigen-level spacing $s_{2j+1}^{o-e}=E_{2j+1}-E_{2j}$ [41, 11]. In the extended phase, the spectrum is nearly doubly degenerate, leading to $s_{2j+1}^{o-e}\sim 0$, so that a noticeable gap exists between $s_{2j}^{e-o}$ and $s_{2j+1}^{o-e}$. In the localized phase, such degeneracies vanish, and the numerically calculated gap sharply decreases in magnitude. In the critical phase, both $s_{2j}^{e-o}$ and $s_{2j+1}^{o-e}$ exhibit strongly scattered distributions, indicative of its multifractal properties [11]. These expected features are clearly visible in Fig.1(d), with the phase boundaries consistent with those in Fig. 1(a)(c).
@@
Figure 1: Fractal dimension FD of $\hat{H}_{p}(a,4-a)$ for (a) $d=2$, and (b) $d=10$. (c) The fractal dimension as a function of the eigenstate index $j/N$, with fixed $a=3$ [the vertical black dashed line in (a)], for $N=2584$ (red dots) and $N=46368$ (black dots), respectively. (d) The even-odd eigen-level spacings $s_{2j}^{e-o}$ (black dots) and $s_{2j+1}^{o-e}$ (red dots) as functions of $j/N$ for $N=28657$. The vertical blue dashed lines in (c) and (d) indicate the positions of the mobility edge, separating four distinct regions corresponding to the localized (I, IV), critical (II), and extended (III) phases. We take $N=2584$ in (a)(b), and impose the periodic boundary condition for all calculations.
@@
In Fig. 2(a)(b), we show the fractal dimensions of the eigenstates of $\hat{H}(a,+\infty)$ and $\hat{H}(+\infty,a)$, respectively. While $\hat{H}(a,+\infty)$ possesses only extended and critical phases, its dual $\hat{H}(+\infty,a)$ supports localized and critical phases. In either case, the phase diagram is symmetric with respect to $j/N=1/2$, which originates from a spectral symmetry: for each eigenvalue $E$, its opposite eigenvalue $-E$ also exists in the spectrum, with both eigenstates featuring the same spatial localization since they differ only by a sign reversal on odd (or even) lattice sites [77]. Furthermore, regions with mobility edges that are constant in $j/N=P_{1,2,3,4}$ can be identified, with $P_{1}=1-\tau$, $P_{2}=20\tau-12$, $P_{3}=7\tau-4$, and $P_{4}=2\tau-1$, all independent of the system size. Note that, although the positions of these mobility edges remain unchanged in terms of the eigenstate index $j/N$ within each region in Fig. 2, their corresponding eigenenergies vary with $a$. These mobility edges separate the extended (red) or localized (blue) phases from the critical phase (green). To better identify these phases, we compute the standard deviation of the fractal dimension [20], defined as $\sigma^{2}(FD)=\langle(FD)^{2}\rangle-\langle FD\rangle^{2}$, where the average is taken over eigenstates in a specified interval. Figure 2(c) shows $\sigma^{2}(FD)$ as a function of $a$ for both $\hat{H}(a,+\infty)$ (black dots) and $\hat{H}(+\infty,a)$ (red dots), with the average taken over the interval $j/N\in[0,P_{4}]$. As expected, $\sigma^{2}(FD)$ exhibits distinct behaviors in different phases: it tends toward zero in the localized region, remains finite in the extended region, and takes intermediate values in the critical region.
@@
$\gamma(E)=\lim_{N\rightarrow\infty}\frac{1}{N}\text{ln}||\Pi_{n=1}^{N}T_{n}||,$ (12)
where |||||$\cdot$||| | ⋅ | | denotes the matrix norm, and the transfer matrix $T_{n}$ is given by
@@
with $t_{n}=\sum_{s=1}^{d}\frac{\cos[\tau s\pi(2n+1)]}{s^{a}}$. Here the calculation is feasible since the transfer matrices are easily accessible for $\hat{H}_{p}(a,+\infty)$ and $\hat{H}_{p}(+\infty,a)$, which is not generally the case for $\hat{H}_{p}(a,b)$. In Fig. 2(d), we show the calculated LE as a function $a$ and $j/N$, the results match excellently with the phase diagram in Fig. 2(b). Specifically, for critical states, the LE vanishes ($\gamma=0$), while for localized states, it remains finite [19].
Figure 2: Fractal dimensions for different eigenstates with $d=2$ for (a) $\hat{H}_{p}(a,+\infty)$ and (b) $\hat{H}_{p}(+\infty,a)$. Four characteristic regions with step-wise mobility edges in $j/N$ can be identified, labeled by $j/N=P_{1,2,3,4}$. (c) The standard deviation of FD, $\sigma^{2}(FD)$, as a function of $a$ for $\hat{H}(a,+\infty)$ (black dots) and $\hat{H}(+\infty,a)$ (red dots), with the average taken over the eigenstates in the interval $j/N\in[0,P_{4}]$. (d) The Lyapunov exponent as a function of $a$ and $j/N$. For all calculations, we take $N=2584$, and impose the periodic boundary condition.
@@
Assuming a constant $\Delta_{n}^{2}+\Omega_{n}^{2}$, we perform the rotating-wave approximation in an appropriate rotating frame, arriving at the Hamiltonian [80, 77]
$\hat{H}_{I}=$ (14)
$\sum_{m\neq n}J_{\xi_{+},mn}\hat{b}_{\xi_{+},m}^{\dagger}\hat{b}_{\xi_{+},n}+J_{\xi_{-},mn}\hat{b}_{\xi_{-},m}^{\dagger}\hat{b}_{\xi_{-},n},$
where $J_{\xi_{+},mn}$ and $J_{\xi_{-},mn}$ are the corresponding coupling coefficients, and the creation operators are $\hat{b}_{\xi_{+},n}^{\dagger}=|\xi_{+}\rangle_{n}\langle 0|_{n}$ and $\hat{b}_{\xi_{-},n}^{\dagger}=|\xi_{-}\rangle_{n}\langle 0|_{n}$.
@@
Further requiring that the energy splittings between the microwave-dressed basis be much larger than the dipole interaction energy [77], we focus on the case where only a single excited state $|\xi_{+}\rangle$ is present in the system. Then, the Hamiltonian simplifies to
$\hat{H}_{I}=\sum_{m\neq n}J_{\xi_{+},mn}\hat{b}_{\xi_{+},m}^{\dagger}\hat{b}_{\xi_{+},n},$ (15)
where the coupling coefficients are
$J_{\xi_{+},mn}=\frac{1}{{|m-n|}^{3}}(\cos\frac{\alpha_{m}-\alpha_{n}}{2}Y_{20}+$ (16)
$\sqrt{6}\cos\frac{\alpha_{m}}{2}\sin\frac{\alpha_{n}}{2}Y_{2-2}+\sqrt{6}\cos\frac{\alpha_{n}}{2}\sin\frac{\alpha_{m}}{2}Y_{22}).$
Here $\cos\alpha_{n}=\frac{\Delta_{n}}{\sqrt{\Delta_{n}^{2}+\Omega_{n}^{2}}}$, $\sin\alpha_{n}=\frac{\Omega_{n}}{\sqrt{\Delta_{n}^{2}+\Omega_{n}^{2}}}$, and $\{Y_{20},Y_{2\pm 2}\}$ are the spherical harmonics. We set $Y_{20}=0$ and $Y_{2-2}=Y_{22}$ by adjusting the relative orientation between the Rydberg-atom array and the spherical coordinate. By further setting appropriate values of $\Delta_{n}$ and $\Omega_{n}$, such that $\alpha_{n}=2\pi\tau n+\frac{\pi}{2}$, the Hamiltonian reduces to
$\ \hat{H}_{\xi_{\pm}}=A\sum_{m\neq n}\frac{\cos[\pi\tau(m+n)]}{|m-n|^{3}}\hat{b}_{\xi_{+},m}^{\dagger}\hat{b}_{\xi_{+},n},$ (17)
which recovers Eq. (10). Here the pre-factor $A$ is proportional to the dipolar interaction energy between adjacent atoms.

### v3

# Family of one-dimensional self-dual quasicrystals with critical phases

@@ Abstract
We propose a general framework for constructing self-dual one-dimensional quasiperiodic lattice models with arbitrary-range hoppings and multifractal behaviors. Our framework generates a broad spectrum of one-dimensional quasicrystals, ranging from the off-diagonal Aubry-André-Harper models on one end, to those featuring long-range hoppings with varied quasiperiodic modulations on another. Focusing on models with off-diagonal quasiperiodic hoppings with power-law decay, we exploit the fact that, when the self-dual condition is satisfied, the system must be in the critical state with multifractal properties. This enables the engineering of models with competing extended, critical, and localized phases, with richly structured mobility edges separating them. As an outstanding example, we show that a limiting case of our family of self-dual quasicrystals can be implemented using Rydberg-atom arrays. Our work offers a systematic route toward critical phases from self-duality considerations, and would facilitate the experimental simulation of these exotic states.
@@
Introduction.— Anderson localization is a fundamental quantum phenomenon in which disorder leads to the exponential localization of eigenstates, thereby inhibiting wave diffusion [1, 2, 3, 4, 5]. The associated location-delocalization transition generally depends on the spatial dimensions, for instance, an infinitesimally weak disorder is sufficient to induce Anderson localization in one dimension [6, 7, 8]. But Anderson localization also arises in quasiperiodic models, where, in a one dimensional quasicrystal for instance, the localization only occurs beyond a critical threshold of the quasiperiodic potential. This is exemplified by the well-known Aubry-André-Harper (AAH) model [9], characterized by the eigen equation $t\left(\psi_{n+1}+\psi_{n-1}\right)+V\cos(2\pi\tau n)\psi_{n}=E\psi_{n}$, where $\psi_{n}$ is the wave function component on site $n$, $t$ is the nearest-neighbor hopping strength, $V$ is the amplitude of the quasiperiodic potential, $\tau$ is an irrational number, and $E$ is the eigenenergy. Crucially, the AAH model exhibits self-duality, such that it Fourier-transforms onto itself but with switched parameters $V\leftrightarrow 2t$. The self-dual point $V=2t$ thus marks a sharp transition between the extended (for $V<2t$) and localized (for $V>2t$) phases. Remarkably, at the self-dual point, the system becomes critically localized [11, 12, 13, 14, 15, 16, 10, 17, 18, 19, 20].
@@
Further engineering of the AAH model can have two important consequences. First, with the addition of quasiperiodic hopping coefficients, the critical point can be expanded into a critical phase occupying a broad region on the phase diagram. These critical states exhibit unique spectral statistics [21, 22, 23], multifractal properties [24, 25, 26], unconventional dynamics [27, 28, 29], and superconductivity [30, 31]. Recent studies have shown that when interactions are introduced into systems supporting single-particle critical states, the resulting models can host many-body critical phases, representing a dynamical regime that is neither ergodic nor many-body localized.  [32, 33, 34]. Second, by introducing long-range hopping terms [35] or fine-tuned on-site quasiperiodic potentials [36], mobility edges emerge in the eigenspectum, where critical energies separate extended and localized states [3, 5, 35, 36, 38, 33, 37, 39]. Tunable mobility edges suggest energy- or parameter-dependent transport, which are highly desirable for device design. These possibilities have thus sparked widespread interest, including theoretical studies on periodically (or quasiperiodically) driven systems [40, 41, 42, 43, 44, 45] and non-Hermitian quasicrystals [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], as well as experimental implementations using ultracold atoms [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], photonic crystals [68, 69, 70, 71, 72, 73, 74], optical cavities [75, 76], and superconducting circuits [77, 11]. Despite these activities, the general mechanisms responsible for the emergence of critical states remain an open question, and the coexistence (in the same phase diagram) of extended, critical, and localized phases has so far been observed only in a handful of highly fine-tuned models [78, 79, 80, 17]. While it is tempting to address these issues from the perspective of self-duality, only a limited number of quasiperiodic models are known to be self-dual [19, 35, 36, 78].
@@
In this work, we propose a general framework for constructing self-dual one-dimensional quasiperiodic models with with arbitrary hopping modulations that give rise to a variety of critical behaviors. Similar to the AAH example, with our scheme, a model constructed with the parameters $(a,b)$ is dual to the same model but parametrized by $(b,a)$. This naturally leads to the self-dual condition $a=b$, where the system must be in the critical state, and remains so in its vicinity. The family of quasicrystals so generated spans a wide spectrum, with the (generalized) AAH models on one end, and those with long-range hoppings having exponential or power-law decay on another. We focus on models exhibiting off-diagonal quasiperiodic hoppings with power-law decay, and show that the overall phase diagram is symmetric with respect to the self-dual condition $a=b$, featuring rich mobility edges in between the extended, critical, and localized phases. We then propose an experimental-simulation scheme using Rydberg-atom arrays. Our work provides a general scheme for engineering self-dual quasicrystals, as well as a route toward critical phases from self-duality considerations.
@@
where the irrational number $\tau=({\sqrt{5}-1})/{2}$. Note that our proposal is still applicable when choosing a different irrational number. Performing the dual transformation
@@
where $N$ represents the total number of sites, the Hamiltonian in the dual space reads
@@
Here the derivation of Hamiltonian (4) only holds for an irrational $\tau$ [81]. Equations (1) and (4) establish that $\hat{H}(a,b)$ and $\hat{H}(b,a)$ are dual to each other, such that $a=b$ defines the self-dual condition. It follows that, if the $j$th eigenstate of $\hat{H}(a,b)$, denoted as $\psi^{(j)}$, is localized, then the $j$th eigenstate of $\hat{H}(b,a)$, denoted as $\tilde{\psi}^{(j)}$, must be extended, and vice versa. Furthermore, if $\psi^{(j)}$ is critical, then $\tilde{\psi}^{(j)}$ is also critical, such that all eigenstates of $\hat{H}(a,a)$ are critical, being at the self-dual point $a=b$. Note that, in the following, we index the eigenstates according to their eigenenergies in ascending order.
Importantly, Hamiltonian $\hat{H}(a,b)$ is determined by the functions $\{f_{s}(a)\}_{s=1,2,...,+\infty}$, which are arbitrary, and subject only to the constraint that the summation in Eq. (2) converge. Such a flexibility allows us to construct a broad class of self-dual quasicrystals with multifractal properties [81].
@@
Quasicrystals with power-law hoppings.— Inspired by the recent progress in Rydberg-atom arrays with long-range interactions, we now focus on a class of quasiperiodic models with power-law hopping terms. Specifically, we consider $f_{s}(a)=\frac{1}{s^{a}}$ for $s\leq d$, and $f_{s}(a)=0$ for $s>d$. In this case, we have
@@
where $d$ denotes the maximum hopping range, while $a$ and $b$ respectively characterize the exponent of the power-law decay in hopping, and the amplitude modulation. Note that for $a,b\to\infty$, Hamiltonian (8) reduces to the off-diagonal AAH model (6).
To corroborate the self-duality and localization properties of the model, we study the fractal dimension (FD) of $\hat{H}_{p}(a,b)$. The FD is given by
@@
As a convenient sampling of the parameter space, we impose the constraint $a+b=4$, such that the self-dual condition is satisfied for $a=2$. In Fig. 1(a)(b), we show the computed FD as functions of $a$ and the eigenstate index $j/N$ for different hopping ranges $d$. The resulting phase diagram exhibits a rich structure, featuring all three phases: the extended (red), the critical (green), and the localized (blue). Consistent with the self-dual properties of the model, the phase diagram is symmetric: regions of the extended (localized) phase to the left of $a=2$ are symmetric to those of the localized (extended) phase to the right, while regions of the critical phase are symmetric with respect to $a=2$. With increasing hopping range $d$, the stability region of the critical phase shrinks and becomes fragmented. Note that hoppings with power-law decay are not essential for our scheme, other forms of decay (exponential for instance) can lead to qualitatively similar phase diagrams [81].
@@
In Fig. 1(c), we show the FD as a function of the eigenstate index $j/N$ for different system sizes. In the localized regions (marked I and IV), FD tends toward zero with increasing system size, while in the extended region (marked III), FD increases toward unity. In contrast, in the critical region (marked II), FD remains between 0 and 1, showing minimal dependence on the system size.
@@
Alternatively, the three phases can be distinguished by analyzing the level statistics. Denoting the ordered eigenenergies as $\{E_{j}\}_{j=1,2,...N}$, we define the even-odd eigen-level spacing as $s_{j}^{e-o}=E_{j}-E_{j-1}$ for even $j$, and $s_{j}^{o-e}=E_{j}-E_{j-1}$ for odd $j$ [44, 10]. In the extended phase, the spectrum is nearly doubly degenerate, leading to $s_{j}^{o-e}\sim 0$, so that a noticeable gap exists between $s_{j}^{e-o}$ and $s_{j}^{o-e}$. In the localized phase, such degeneracies vanish, and the numerically calculated gap sharply decreases in magnitude. In the critical phase, both $s_{j}^{e-o}$ and $s_{j}^{o-e}$ exhibit strongly scattered distributions, indicative of its multifractal properties [10]. These expected features are clearly visible in Fig.1(d), with the phase boundaries consistent with those in Figs. 1(a) and 1(c).
@@
Figure 1: Fractal dimension FD of $\hat{H}_{p}(a,4-a)$ for (a) $d=2$, and (b) $d=10$. (c) The FD as a function of the eigenstate index $j/N$, with fixed $a=3$ [the vertical black dashed line in (a)], for $N=2584$ (red dots) and N=46 368N=46\,368 (black dots), respectively. (d) The even-odd eigen-level spacings $s_{j}^{e-o}$ for even $j$ (black dots) and $s_{j}^{o-e}$ for odd $j$ (red dots), both as functions of $j/N$ for N=28 657N=28\,657. The vertical blue dashed lines in (c) and (d) indicate the positions of the mobility edge, separating four distinct regions corresponding to the localized (I, IV), critical (II), and extended (III) phases. We take $N=2584$ in (a)(b), and impose the periodic boundary condition for all calculations. For numerical calculations, we approximate the golden ratio $\tau$ as the ratio of two adjacent $u$-Fibonacci numbers, with $\tau=1597/2584$ here [55, 81].
Figure 2: The FDs for different eigenstates with $d=2$ for (a) $\hat{H}_{p}(a,+\infty)$ and (b) $\hat{H}_{p}(+\infty,a)$. Four characteristic regions with step-wise mobility edges in $j/N$ can be identified, labeled by $j/N=P_{1,2,3,4}$. (c) MFDs as functions of $1/\log_{10}N$ for eigenstates in different spectral regions for $a=3$ in (a)(b). Red squares and blue circles correspond to the extended [in (a)] and localized phases [in (b)], respectively. Green triangles and stars indicate the MFDs for the critical states in (a) and (b), respectively. Solid lines denote results of linear fit. (d) The Lyapunov exponent as a function of $a$ and $j/N$. We take $N=2584$ in (a)(b)(d), and impose the periodic boundary condition for all calculations.
Figure 3: (a) Level scheme of the microwave-dressed Rydberg states of the $n$th atom in the one-dimensional array. Here $2\Delta_{n}$ is the site-dependent energy splitting between the Rydberg states $|+\rangle_{n}$ and $|-\rangle_{n}$, $\Omega_{n}$ is the Rabi frequency of the coupling between $|+\rangle_{n}$ and $|-\rangle_{n}$. (b) Level scheme for the microwave-dressed basis of the $n$th atom. (c)(d) Fractal dimension FD of different eigenstates with varying interaction (hopping) range $d$ for (c) $H(3,+\infty)$ and (d) $H(+\infty,3)$, respectively. The two mobility edges are located at $j/N=2-2\tau$ and $2\tau-1$, respectively. For all calculations, the system size is $N=2584$ and we impose the periodic boundary condition.
@@
In Figs. 2(a) and 2(b), we show the FDs of the eigenstates of $\hat{H}(a,+\infty)$ and $\hat{H}(+\infty,a)$, respectively. While $\hat{H}(a,+\infty)$ possesses only extended and critical phases, its dual $\hat{H}(+\infty,a)$ supports localized and critical phases. In either case, the phase diagram is symmetric with respect to $j/N=1/2$, which originates from a spectral symmetry: for each eigenvalue $E$, its opposite eigenvalue E-E also exists in the spectrum, with both eigenstates featuring the same spatial localization since they differ only by a sign reversal on odd (or even) lattice sites [81]. Furthermore, regions with mobility edges that are constant in $j/N=P_{1,2,3,4}$ can be identified, with $P_{1}=1-\tau$, $P_{2}=20\tau-12$, $P_{3}=7\tau-4$, and $P_{4}=2\tau-1$, all independent of the system size. Here the expressions of $P_{n}$ are determined through numerical analysis, with their forms sensitive to the choice of $\tau$ [81]. Further, although the positions of these mobility edges remain unchanged in terms of the eigenstate index $j/N$ within each region in Fig. 2, their corresponding eigenenergies vary with $a$. These mobility edges separate the extended (red) or localized (blue) phases from the critical phase (green). For a fermionic system, the step-wise mobility edges in $j/N$ suggest a density-dependent quantum transport that can be robustly tuned through the parameter $a$.
To better confirm the nature of different phases, we perform a finite-size scaling of the mean fractal dimension (MFD). Take $a=3$ as an example, the MFD in the critical region is defined by averaging the FD of all eigenstates with $j/N\in[P_{4},1-P_{4}]$. The MFD in the extended (localized) region is averaged outside this interval. Figure 2(c) shows the MFDs as functions of $1/\log_{10}N$ for different phases of $\hat{H}_{p}(3,+\infty)$ and $\hat{H}_{p}(+\infty,3)$, respectively. Crucially, in the thermodynamic limit $1/\log_{10}N\to 0$, the MFDs in the critical regions extrapolate to finite values in $(0,1)$, consistent with the multifractal nature of the critical states and confirming the critical nature of the model away from the self-dual point.
@@
[…]
where $||\cdot||$ denotes the matrix norm, corresponding to taking the largest absolute value among the eigenvalues. The transfer matrix $T_{n}$ is given by
@@
with $t_{n}=\sum_{s=1}^{d}\frac{\cos[\tau s\pi(2n+1)]}{s^{a}}$. Here the calculation is feasible since the transfer matrices are easily accessible for $\hat{H}_{p}(a,+\infty)$ and $\hat{H}_{p}(+\infty,a)$, which is not generally the case for $\hat{H}_{p}(a,b)$. In Fig. 2(d), we show the calculated LE as a function of $a$ and $j/N$, the results match excellently with the phase diagram in Fig. 2(b). Specifically, for critical states, the LE vanishes ($\gamma=0$), while for localized states, it remains finite [18].
@@
Performing the rotating-wave approximation in an appropriate rotating frame, and setting appropriate values of $\Delta_{n}$ and $\Omega_{n}$, we reproduce Hamiltonian (10) as [84, 81]
[…]
@@
where $\hat{b}_{\xi_{+},n}^{\dagger}=|\xi_{+}\rangle_{n}\langle 0|_{n}$. For the derivation, we require that the energy splittings between the microwave-dressed basis be much larger than the dipole interaction energy [81], and focus on the case where only a single excited state $|\xi_{+}\rangle$ is present. Here the pre-factor $A$ is proportional to the dipolar interaction energy between adjacent atoms. Under typical experimental conditions, it suffices to take the hopping range $d=2$ or $3$ [85, 86, 87].